# Explicit Visual Prompt Instruction Tuning

### v1

@@ Abstract
Large Multimodal Models (LMMs) have made significant breakthroughs with the advancement of instruction tuning. However, while existing models can understand images and videos at a holistic level, they still struggle with instance-level understanding that requires a more nuanced comprehension and alignment. Instance-level understanding is crucial, as it focuses on the specific elements that we are most interested in. Excitingly, existing works find that the state-of-the-art LMMs exhibit strong instance understanding capabilities when provided with explicit visual cues. Motivated by this, we introduce an automated annotation pipeline assisted by GPT-4o to extract instance-level information from images and videos through explicit visual prompting for instance guidance. Building upon this pipeline, we proposed Inst-IT, a solution to enhance LMMs in Instance understanding via explicit visual prompt Instruction Tuning. Inst-IT consists of a benchmark to diagnose multimodal instance-level understanding, a large-scale instruction-tuning dataset, and a continuous instruction-tuning training paradigm to effectively enhance spatial-temporal instance understanding capabilities of existing LMMs. Experimental results show that, with the boost of Inst-IT, our models not only achieve outstanding performance on Inst-IT Bench but also demonstrate significant improvements across various generic image and video understanding benchmarks. This highlights that our dataset not only boosts instance-level understanding but also strengthens the overall capabilities of generic image and video comprehension.
Equal contributions. Corresponding author.
Figure 1: Current LMMs struggle with instance-level understanding, failing to capture the nuanced details about specific instances. To address this, we create a large-scale instance-specific instruction tuning dataset and train a multimodal model on it. Compared to existing models, our model show better performance in instance-level understanding.
@@ 1 Introduction
Recently, Large Multimodal Models (LMMs) have seen remarkable advancements. A key breakthrough is visual instruction tuning [39, 16], enabling models to follow any type of user instructions. This paves the way to building general-purpose multimodal assistants capable of handling a wide range of real-world tasks [30]. Inspired by this initial work, numerous follow-up studies have emerged in both image-language [40, 54, 99, 10, 12, 73] and video-language [46, 37, 90, 33, 72, 76, 74] domains. However, although these models can understand the entire images or videos, they still struggle to comprehend instance-specific content, as illustrated in Fig. 1. This hinders them from a fine-grained understanding of visual inputs.
@@
Instance-level understanding refers to comprehending the attributes of specific instances within an image or video, as well as the relationships and interactions between them. This requires models to exhibit nuanced comprehension and fine-grained alignment. Instance understanding has been a long-standing pursuit of the computer vision community with extensive efforts devoted to object detection [65, 63, 52], instance segmentation [66, 53] and object tracking [18], etc. This capability is essential for real-world applications, where users pay more attention to the instances that they are interested in. In the era of LMMs, although there have been some attempts in exploring multimodal instance understanding [22, 5, 94, 98, 95], they are primarily limited in the image domain, leaving the videos under-explored. Compared to images, understanding instances in videos is considerably more challenging, as it requires not only capturing their spatial information but also their temporal dynamics. Driven by this, we aim to advance the multimodal instance-level understanding in both images and videos. To this end, we focus on three key aspects of the effort: the evaluation benchmark, the instruction-tuning dataset, and the training recipe.
We begin by introducing an automated data annotation pipeline designed to provide detailed annotations for images and videos, with a particular emphasis on individual instances. Specifically, this pipeline first highlights the instances of interest in the images and videos using instance-level visual prompts [81]. Then, GPT-4o [55] is leveraged to generate fine-grained and multi-level annotations that covering several aspects, including: instance-level captions, frame-level captions, temporal changes descriptions, video-level captions, and open-ended question-answer pairs.
With the help of this pipeline, we first carefully design the Inst-IT Bench, a benchmark for diagnosing multimodal models in instance-level understanding, which includes 1,000 image test samples and 1,000 video test samples. We conduct extensive evaluations on Inst-IT Bench, and find that existing open-source models show limited performance in understanding individual instances. To address this gap, we apply the propose data annotation pipeline to perform fine-grained annotations on nearly 21k videos and 51k images, resulting in Inst-IT Dataset, which contains a total of 207k frame-level annotations, 21k video-level annotations, and 335k open-ended question answer pairs. To the best of our knowledge, this is the first instruction-tuning dataset that provides fine-grained annotations focused on instance of interest in both images and videos. Based on Inst-IT Dataset, we propose a continuous instruction-tuning recipe that effectively combines our instance understanding datasets with other general instruction-tuning data. By incorporating this small amount of data, the enhanced models demonstrate strong performance not only on Inst-IT Bench, but also on several widely used image and video benchmarks, such as AI2D [25] and Egoschema [47].
Our contributions are three-fold:
@@
    An instance-centric understanding benchmark. We introduce Inst-IT Bench, a benchmark specifically designed to evaluate instance-level understanding in multimodal models, encompassing both image and video understanding, and perform extensive evaluations on it.
@@
    An instance-grounded instruction-tuning dataset. We develop Inst-IT Dataset, the first instruction-tuning dataset that includes both images and videos, featuring explicit instance-level visual prompts. It is built through a GPT-4o-assisted automated annotation pipeline.
@@
    An instance-enhanced Large Multimodal Model. We further integrate Inst-IT Dataset into the instruction tuning of LMMs. We propose a continuous instruction tuning approach to effectively enhance spatial-temporal instance understanding while consistently improving general comprehension.
@@ 2 Related Work
Large Multimodal Models. Recently, significant progress has been witnessed in large multimodal models (LMMs). BLIP-2 [32] and Flamingo [2] leverage visual resamplers to integrate image features as language inputs by extracting a fixed number of visual tokens. In a parallel line of research, LLaVA [39] and its follow-upds [54, 40, 29, 97, 38, 51, 91] have achieved remarkable success by connecting vision and language through a simple projection module. This approach has simplified the pipeline, even outperforming previous models with less training effort. Additionally, researchers are extending LMMs’ capabilities to temporal understanding by incorporating multi-frame inputs [37, 74] or explicit temporal modules [36, 23], bringing these models closer to applications in real-world scenarios. However, we observe that existing LMMs struggle with instance-level understanding and often fail to accurately follow instructions to ground specific instances. We emphasize the importance of instance understanding and enhance it through instruction fine-tuning with explicit visual prompts, which also enhances general understanding.
@@
Multimodal Datasets and Benchmarks. With the rapid progress in LMMs, numerous instruction tuning datasets have been developed. LLaVA-Instruct [39] is one of the first to use GPT-4, leveraging object categories, bounding boxes, and image-level captions to generate diverse visual instruction tuning data. Building on this, follow-up studies have focused on enhancing large language and multi-modal models to generate synthetic data [10, 73, 8] or improve the annotation pipeline [11, 97]. Simultaneously, various multimodal benchmarks have been proposed to evaluate LMMs across different aspects, such as comprehensive understanding [28], OCR and text comprehension [48, 49, 50, 70], temporal understanding [21, 47, 77, 6, 42], hallucination [34], and instruction-following [60]. However, they focus more on image or video-level understanding and lack fine-grained emphasis on specific instances. In this work, we emphasize the importance of instance understanding in both images and videos. Specifically, we propose the Inst-IT Bench to evaluate the instance understanding of LMMs and create the Inst-IT Dataset, providing detailed instance-level annotations to enhance instance understanding capabilities.
@@
Multimodal Instance Understanding. Understanding individual instances in visual inputs is a central focus in computer vision, with key tasks such as object detection [65, 63], instance segmentation [66, 27], and object tracking [18, 45, 84]. In the era of large multimodal models, instance-level understanding has also gained increasing attention. SPEC [56], ARO[88] and Winoground [69], evaluates the ability of vision-language models to comprehend fine-grained visual-linguistic concepts, revealing that models like CLIP [61] struggle to grasp attributes of specific instances or the relationships between them. To address this, works like KOSMOS-2 [57] and Shikra [9] encode instance information in textual form. In parallel, another line of research, including SoM-LLaVA [79], RegionGPT [22], GPT4ROI [94], MG-LLaVA [98], OMG-LLaVA [95], and ViP-LLaVA [5], explores the use of visual prompting to guide models in focusing on specific instances. Among these, SoM-LLaVA [79] is most closely related to our work. It uses SoM [81] to prompt images and asks models to list the instances present, finding that instance-level understanding enhances global understanding. However, SoM-LLaVA [79] focuses solely on the image domain and does not address the more challenging video domain. In contrast, we focus on both images and videos, aiming to advance multimodal models in understanding the spatiotemporal changes of individual instances.
Figure 2: The automated data generation pipeline. We process the video frames sequentially. At each timestamp $t$, GPT-4o is prompted to create a frame-level annotation $Y_{t}^{f}$ based on the current frame $X_{t}$ and the previous frame $X_{t\text{-}1}$. Then, all the frame-level annotations are aggregated to produce a video-level description $Y^{vid}$ and a set of open-ended question-answer pairs $Y^{qa}$.
Figure 3: Visualization of an data example from our Inst-IT Dataset. For each video, we provide (a) frame-level annotations, (b) a video-level description, and (c) open-ended question-answer pairs. Each frame-level annotation consists of three parts: captions for individual instances, a caption for the entire scene, and captions describing the temporal changes involving specific instances. A key feature of our dataset is its emphasis on instances of interest, including their state in each frame, how they change between frames, and questions and answers focused on their specific details throughout the video. The contours of instances in this example are deliberately highlighted for better visualization. A complete data example can be found in Sec. C.2.
@@ 3 Inst-IT
High-quality instruction following data are essential for training effective multimodal models. To address the shortage of instance-grounded data, we propose an automated pipeline to generate detailed annotations for both images and videos, with a special emphasis on Instances of Interest. We begin by collecting a diverse set of images and videos, construct visual prompts and leverage GPT-4o to generate instruction-following data based on these visual prompted inputs (Sec. 3.1). Building upon this pipeline, we carefully design a benchmark to assess models in instance-specific understanding (Sec. 3.2). Furthermore, we build a large-scale vision language dataset (Sec. 3.3) and proposed a continuous instruction-tuning pipeline (Sec. 3.4) to enhance the instance understanding capabilities of LMMs.
@@ 3.1 Instance-centric Annotation Pipeline
Overview. We propose an automated annotation pipeline designed to annotate videos with a particular focus on individual instances. As illustrated in Fig. 2, the overall process involves annotating each frame sequentially, combining these frame-level annotations into a comprehensive video-level description, and generating a set of open-ended questions about the video content.
Visual Prompting. To generate more fine-grained annotations focused on specific instances, we augment the images and videos with visual prompts to highlight the instances of interest. Specifically, we use the Set-of-Marks (SoM) visual prompting [81], which overlays a numerical ID directly on each instance. We find this method highly effective in guiding GPT-4o to provide annotations focused on individual instances. For more details on SoM prompting, please refer to Appendix A.
Frame-level Annotation. We annotate each frame of the video sequentially. At each timestamp $t$, we provide GPT-4o with the current frame $X_{t}$, the previous frame $X_{t\text{-}1}$, and a tailored task prompt $P^{f}$. GPT-4o then produces a frame-level annotation $Y_{t}^{f}\text{=}(y_{t}^{ins},y_{t}^{img},y_{t}^{dif})$ across three aspects, where $y_{t}^{ins}$ represents the captions for individual instances, $y_{t}^{img}$ is caption for the entire image, and $y_{t}^{dif}$ describes the temporal differences from the previous frame:
$Y_{t}^{f}=\texttt{GPT}(P^{f},X_{t},X_{t\text{-}1})$ (1)
Video-level Summary. After obtaining annotations for each frame, we employ GPT-4o to aggregate them into $Y^{vid}$, a comprehensive description for the entire video:
$Y^{vid}=\texttt{GPT}(P^{vid},[Y_{1},Y_{2},\cdots,Y_{N}]),$ (2)
@@
Open-Ended Question Answer Pairs. Based on the frame-by-frame annotations, we also prompt GPT-4o with the task prompt $P^{qa}$ to create $M$ open-ended QA pairs $Y^{qa}\text{=}\{(q_{i},a_{i})\}_{i=1}^{M}$ focusing on specific instances:
$Y^{qa}=\texttt{GPT}(P^{qa},[Y_{1},Y_{2},\cdots,Y_{N}])$ (3)
Putting It Altogether. Through the steps outlined above, we can annotate the video content at multiple levels of granularity, highlighting the elements related to specific instances. Specifically, as illustrated in Fig. 3, each video is accompanied by the following three types of annotations:
  • $N$ frame-level annotations, each encompassing a detailed description of individual instances, the entire image, and temporal changes.
@@
  • $M$ open-ended question-answer pairs that focused on specific instances or the inter-instance relationships.
Additional information about the design of each task prompts, as well as the organization of the input for GPT-4o is provided in Appendix A.

### v2

# Inst-IT: Boosting Instance Understanding via Explicit Visual Prompt Instruction Tuning

@@ Abstract
Large Multimodal Models (LMMs) have made significant breakthroughs with the advancement of instruction tuning. However, while existing models can understand images and videos at a holistic level, they still struggle with instance-level understanding that requires a more fine-grained comprehension and alignment. Instance-level understanding is crucial for LMMs, as it focuses on the specific elements that we are most interested in. Excitingly, existing works find that the state-of-the-art LMMs exhibit strong instance understanding capabilities when provided with explicit visual cues. Motivated by this, we proposed Inst-IT, a solution to enhance LMMs in Instance understanding via explicit visual prompt Instruction Tuning for instance guidance. Inst-IT consists of a benchmark to diagnose multimodal instance-level understanding, a large-scale instruction-tuning dataset, and a continuous instruction-tuning training paradigm to effectively enhance spatial-temporal instance understanding capabilities of existing LMMs. Experimental results show that, enhanced by Inst-IT, our models not only achieve outstanding performance on Inst-IT Bench and other instance understanding benchmarks, but also demonstrate significant improvements across various generic image and video understanding benchmarks. This highlights that our method not only boosts instance-level understanding but also strengthens the overall capabilities of generic image and video comprehension.
[…]
[Website]https://inst-it.github.io footnotetext: Equal Contribution.
Corresponding authors.
@@ 1 Introduction
Recently, Large Multimodal Models (LMMs) have seen remarkable advancements. A key breakthrough is visual instruction tuning [llava, instructblip], enabling models to follow any type of user instructions. This paves the way to building general-purpose multimodal assistants capable of handling a wide range of real-world tasks [li2024multimodal]. Inspired by this initial work, numerous follow-up studies have emerged in both image-language [llava1d5, deepstack, minigpt4, sharegpt4v, internvl] and video-language [video-chatgpt, fu2024objectrelator, videollama, openvclip, Qwen2VL] modeling. However, although they can understand images or videos at a holistic level, they still struggle to comprehend instance-specific content that the users are most interested in, as illustrated in figure˜1 (a).
@@
Instance-level understanding involves comprehending the attributes of individual instances within an image or video, as well as the relationships and interactions between them. This requires models to exhibit nuanced comprehension and fine-grained alignment. Instance understanding has been a long-standing pursuit of the community with extensive efforts devoted to object detection [faster_rcnn, yolo, fu2024cross, meng2023detection], instance segmentation [unet, sam, Meng2023SEGICUT], and object tracking [mot, tan2024xtrack]. This capability is essential for real-world applications, where users pay more attention to the instances that they are interested in. In the era of LMMs, although there have been some attempts in exploring multimodal instance understanding [regiongpt, vipllava, zhang2023gpt4roi, mgllava, omgllava], they are primarily limited in the image domain, leaving the videos under-explored. Compared to images, understanding instances in videos is considerably more challenging, as it requires not only capturing their spatial information but also temporal dynamics. Driven by this, we aim to advance the multimodal instance understanding in both images and videos. To this end, we focus on three aspects: instruction-tuning dataset, evaluation benchmark, and training recipe.
Existing multimodal benchmarks and datasets primarily provide coarse-grained knowledge for images and videos, lacking fine-grained annotations for individual instances. To address this, we introduce an automated pipeline to generate detailed instance-specific annotations. As illustrated in figure˜1 (b), we leverage GPT-4o [openai2024gpt4o] to produce multi-level annotations, including instance-level descriptions, image-level captions, temporal dynamics, video-level summaries, and open-ended question-answer pairs. To fully unleash the capability of GPT-4o for more accurate annotations, we systematically design task prompts and employ set-of-marks visual prompts [som] to highlight instances in the visual inputs. Powered by this pipeline, we construct Inst-IT Dataset, an instance-grounded multimodal dataset comprising 51k images and 21k videos, 207k image-level captions, 135k temporal dynamics descriptions, 21k video-level captions, and 335k open-ended question-answer pairs. Furthermore, we carefully design the Inst-IT Bench to diagnose the instance-level understanding capabilities of LMMs, and perform rigorous manual verification and refinement to ensure its data quality.
Figure 1: (a) LMMs struggle with instance understanding, failing to capture the nuanced details of instances specified in user queries. (b) Our instance-centric data annotation pipeline, providing multi-level annotations for individual instances in images and videos.
Building on Inst-IT Dataset, we propose a continuous instruction tuning recipe that effectively integrates our instance understanding datasets with general instruction-tuning data. We augment images and videos with visual prompts, and convert the fine-grained annotations from Inst-IT Dataset into instruction tuning format, emphasizing the model’s spatiotemporal understanding of individual instances. Experimental results show that our enhanced models achieve strong instance understanding performance not only on Inst-IT Bench, but also demonstrate consistent improvements on other instance understanding benchmarks e.g. RefCOCOg [refcocog] and ViP-Bench [vipllava]. We also investigate the models’ general comprehension capabilities on widely used generic benchmarks. The results reveal significant improvements over the baseline, achieving 4.4% and 13.5% gains on AI2D [kembhavi2016diagram] and ChartQA [masry2022chartqa] image benchmarks, as well as 7.8% and 11.8% improvements on Egoschema [mangalam2023egoschema] and NExT-QA [xiao2021next] video benchmarks, respectively. This highlights the effectiveness of Inst-IT in boosting instance understanding while strengthening general comprehension in both images and videos. Our contributions are three-fold:
@@
    We construct the Inst-IT Dataset, the first instance-grounded instruction-tuning dataset that includes both images and videos, featuring explicit instance-level visual prompts and fine-grained annotations grounded on individual instances.
@@
    We introduce the Inst-IT Bench, a human-verified benchmark specifically designed to evaluate the instance-level understanding capabilities of LMMs on both images and videos.
@@
    We propose a continuous instruction tuning recipe, which leverages our instance-level dataset alongside general data, effectively enhancing models in instance understanding while consistently improving general comprehension in both images and videos.
@@ 2 Related Work
Large multimodal models. Recently, significant progress has been witnessed in LMMs [yin2024survey]. BLIP-2 [li2023blip] and Flamingo [alayrac2022flamingo] leverage visual re-samplers to integrate image features as language inputs by extracting a fixed number of visual tokens. LLaVA [llava] and its follow-ups [llava1d5, li2024llava, lin2024vila, mckinzie2025mm1, zhang2024mm1, deepstack, chen2025comp] achieve remarkable success by connecting vision and language through a simple projection module. Additionally, researchers are extending LMMs’ capabilities to temporal understanding by incorporating multi-frame inputs [videollava, Qwen2VL, llavavideo] or explicit temporal modules [li2025llama, he2024ma] However, existing LMMs struggle with instance-level understanding and often fail to accurately follow instructions to ground specific instances. We emphasize the importance of instance understanding and enhance it through instruction fine-tuning with explicit visual prompts.
@@
Multimodal datasets and benchmarks. With the rapid progress in LMMs, numerous instruction-tuning datasets have been developed. LLaVA-Instruct [llava] leverages object categories, bounding boxes, and image-level captions to generate diverse visual instruction tuning data. Follow-up studies use more powerful models to generate synthetic data [sharegpt4v, wang2023see, chen2024allava] and improve the annotation pipeline [li2024densefusion, sharegpt4video, llavavideo]. Simultaneously, various benchmarks are proposed to evaluate LMMs across different aspects [mmesurvey, li2024survey, li2025benchmark], such as comprehensive understanding [li2023seed], OCR [masry2022chartqa, mathew2021docvqa, mathew2022infographicvqa, tito2023hierarchical], temporal understanding [fu2024video, mangalam2023egoschema, xiao2021next, cai2024temporalbench, liu2024tempcompass, liu2024bench], and instruction-following [qian2024mia]. However, they focus more on image or video-level understanding and lack fine-grained emphasis on specific instances. We emphasize the importance of instance understanding in both images and videos, and propose the Inst-IT Bench to evaluate the instance understanding of LMMs and create the Inst-IT Dataset, providing detailed instance-level annotations to enhance instance understanding.
@@
Multimodal instance understanding. Understanding individual instances is a central focus in computer vision community, with key tasks like object detection [faster_rcnn, yolo, chen2025comprehensive], instance segmentation [unet, sam], and object tracking [mot, motsurvey, yilmaz2006object]. In the era of LMMs, instance understanding gains increasing attention. SPEC [spec], ARO[aro], and Winoground [thrush2022winoground] reveal that CLIP [clip] struggle to understand instances. To address this, KOSMOS-2 [Peng2023Kosmos2GM], Ferret [ferret], GLaMM [glamm] and Shikra [chen2023shikra] encode instance information in textual form. In parallel, SoM-LLaVA [somllava], RegionGPT [regiongpt], GPT4ROI [zhang2023gpt4roi], MG-LLaVA [mgllava], OMG-LLaVA [omgllava], and ViP-LLaVA [vipllava], explores the use of visual prompting to guide models in focusing on specific instances. SoM-LLaVA [somllava] and Elysium [wang2024elysium] are closely related to ours. SoM-LLaVA [somllava] asks models to list the instances in images, finding this effective in enhancing model comprehension. However, it is limited to the image domain. Elysium [wang2024elysium] focuses on object understanding in videos but employs relatively simplistic instance annotations. In contrast, we focus on both images and videos and provide multi-level fine-grained annotations for instances, aiming to advance multimodal models in understanding the spatiotemporal dynamics of individual instances.
@@ 3 Inst-IT
To address the scarcity of instance-grounded data, we propose an automated pipeline to generate detailed annotations for both images and videos, with a particular emphasis on instances of interest (Sec.˜3.1). Based on this, we build a large-scale instance-grounded multimodal dataset (Sec.˜3.2), and carefully design an instance-centric evaluation benchmark (Sec.˜3.3). Furthermore, we propose a continuous instruction-tuning recipe (Sec.˜3.4) to enhance LMMs in instance understanding.
@@ 3.1 Instance-centric annotation pipeline
Overview. We propose an automated pipeline to generate annotations grounded on individual instances. As in figure˜1 (b), we annotate each frame sequentially, aggregate frame-level annotations into a comprehensive video-level description, and generate open-ended question-answer pairs.
Visual prompting. Directly processing the raw visual inputs suffers from hallucinations and distraction. To mitigate this issue, we augment the images and videos with visual prompts to highlight the instances. Specifically, we use set-of-marks (SoMs) visual prompt [som], which overlays a numerical ID on each instance. We find this method highly effective in guiding GPT-4o to provide annotations focused on individual instances. For more details, please refer to Sec.˜A.1.
Frame-level annotation. We annotate video frames sequentially. At timestamp $t$, we provide GPT-4o with the current frame $X_{t}$, the previous frame $X_{t\text{-}1}$, and a tailored task prompt $P^{f}$. We then obtain a frame-level annotation $Y_{t}^{f}\text{=}\left(y_{t}^{ins},y_{t}^{img},y_{t}^{dif}\right)$ encompassing three aspects, where $y_{t}^{ins}$ represents the captions for individual instances, $y_{t}^{img}$ is a caption for the entire image, and $y_{t}^{dif}$ describes the temporal differences from the previous frame:
$Y_{t}^{f}=\text{GPT}\left(P^{f},X_{t},X_{t-1}\right).$ (1)
Video-level summary. After obtaining annotations for each frame, we aggregate them into a caption for the entire video $Y^{vid}$, capturing detailed spatiotemporal information of individual instances:
$Y^{vid}=\text{GPT}\left(P^{vid},[Y_{1}^{f},Y_{2}^{f},\cdots,Y_{N}^{f}]\right),$ (2)
@@
Open-ended question-answer pairs. We also prompt GPT-4o with the task prompt $P^{qa}$ to create $M$ open-ended QA pairs $Y^{qa}\text{=}\{(q_{i},a_{i})\}_{i=1}^{M}$ focusing on instance understanding:
$Y^{qa}=\text{GPT}\left(P^{qa},[Y_{1}^{f},Y_{2}^{f},\cdots,Y_{N}^{f}]\right).$ (3)
Following these steps, each video is enriched with multi-granularity annotations that incorporate instance-specific information. As illustrated in figure˜2, these annotations include the following aspects:
  • NN frame-level annotations, each contains detailed descriptions of individual instances, the entire image, and the temporal dynamics between adjacent frames.
@@
  • $M$ open-ended question-answer pairs that focused on individual instances or their relationships.
Additional information about the design of each task prompt is provided in Sec.˜A.2.
Figure 2: Visualization of data structure in Inst-IT Dataset. For each video, we provide (a) NN frame-level annotations, (b) a video-level description, and (c) $M$ open-ended question-answer pairs. A complete example data can be found in Sec.˜C.3.
3.2 Inst-IT Dataset
Instruction tuning plays a crucial role in multimodal training; however, the lack of instance-level datasets hinders the advancement of instance understanding. Using the data annotation pipeline described in Sec.˜3.1, we create a large-scale instruction-tuning dataset, the Inst-IT Dataset. To the best of our knowledge, this is the first instruction fine-tuning dataset that provides multi-level fine-grained annotations centric on individual instances in both images and videos.
Data sources. We utilize five video instance segmentation datasets (BRUST [brust, tao], UVO [uvo], OVIS [ovis], LVVIS [lvvis] and YoutubeVIS-2021 [ytvis]) and two object tracking datasets (BenSMOT [bensmot], VidOR [vidor]) as our video sources, as they provide annotations of instance locations, which is useful in SoM visual prompting [som]. For the image source, we select the SA-1B [sam] dataset due to its diversity and abundance of instance objects. In total, we collect 51k images and 21k videos. More details can be found in Sec.˜C.1.
Statistics. On average, each video includes one video-level annotation, 7.3 frame-level annotations, and 15.6 open-ended QA pairs. Images are regarded as single-frame videos without temporal changes. In total, Inst-IT Dataset includes 21k videos and 51k images, alongside 21k video-level captions, 207k frame-level captions, 836k instance-level captions, 135k temporal descriptions, and 335k open-ended QA pairs. More statistical analyses are provided in Sec.˜C.2.
[…]
Comparison with existing instruction tuning datasets. Tab.˜1 (left) compares Inst-IT Dataset with other datasets. Prior video datasets, e.g. ShareGPT4Video [sharegpt4video] and LLaVA-Video [llavavideo], focus on holistic understanding without instance-level annotations. While VIP-LLaVA [vipllava] offers instance annotations for images, it does not include any video data. In contrast, Inst-IT Dataset encompasses both images and videos with multi-level, fine-grained annotations grounded on individual instances.